OpenStudy (anonymous):

what does 2(small 2) x 5 (small 2) equal??? (exponents)

OpenStudy (anonymous):

Is it: \[2^{2}*5^{2}\]

OpenStudy (anonymous):

If so they want (2*2)*(5*5) 4*25 = 100

OpenStudy (anonymous):

yes it is but they didnt teach me this stuff IM SO CONFUSED!!!

OpenStudy (anonymous):

exponents are a short hand way of writing multiplication. you multiply the base (bottom) number by itself the amount of times the exponents (the small number) tell you to

OpenStudy (anonymous):

So \[4^{3}\] = 4*4*4

OpenStudy (anonymous):

oh so then after u multiply the first set of numbers time the 2nd set?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

can u help me find the answer for this one question please!

OpenStudy (anonymous):

2^2*5^2 =(2*2)*(5*5) =100

OpenStudy (anonymous):

the small value means tha u must multiply the base its self looking that number of small value

OpenStudy (anonymous):

no its use a factor tree to find the prime factorizationof the composite number 50. which answer expresses the number in exponential notation for 84 (multiple chocie)

OpenStudy (anonymous):

choice i mean

OpenStudy (anonymous):

For 50 I would start by factoring out 2's until I can't 2*25 2*5*5 = 2*5^2

OpenStudy (anonymous):

If you do the same for 84: 2*42 2*2*21 2*2*3*7 \[2^{2}*3*7\]

OpenStudy (anonymous):

ok thanks a lot i wouldve never passed this question with out ur help

OpenStudy (anonymous):

remember to hit the "good answer" button on the left so people know your question has been answered :D

OpenStudy (anonymous):

lol ok thanks again